The speed of light in a leaded glass is 1.66x10^7 m/s. What is the index of refraction of this type of glass? 


1
Expert's answer
2021-08-16T08:32:13-0400

The index of refraction

"n=\frac{c}{v}=\frac{3.00*10^8\:\rm m/s}{1.66*10^7\:\rm m/s}=18.1"